The mean score of sixth-formers in Fiji in a national math exam is 75 with a standard deviation of 8.1. A random sample of 100 students is chosen from one school district in Fiji to see if the mean score in that district is lower than the national average. The mean score of these 100 students is 71. Does this indicate that the mean score of sixth-formers in that district is significantly higher than 75 at the 5% level of significance? 


1
Expert's answer
2021-06-08T10:57:30-0400
"S_x=\frac{8.1}{\sqrt{100}}=0.81"

"z=\frac {75-71}{0.81}=4.9"

71+4.9=75.9